Two identical balls of equal masses A & B,arelying on a smooth surface with ball A behind ball B.Ball A hits the ball B(which is at rest) with a velocity v=16 m/s.What should be the minimum value of coefficient of restituion e between A and B so that B just reaches the highest point of inclined plane which is 5m high?Here ball B is near to inclined plane.


Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear sindhuja

initial velocity of A = 16

initial velocity of B = 0

let velocity of A after collision =Va

let velocity of B after collision =Vb

1/2 m Vb2=mg*5

  Vb=10 m/sec

conservation of moment

m*16 +m*0 =m*Va +m*10

Va=6 m/sec

now e =ve,ocity of seperation /velocity of apporoch

        =(10-6) /16

        =1/4

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the

answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin